# A Small Collusion is All You Need

### v1

@@ 4. Corollaries Of The Main Characterization
Consider any $2$-SCP auction $a,p,\beta$. By our characterization, it is also SCP. (Gafni and Yaish, 2025) show that SCP implies Global-SCP, and provide a characterization of Global-SCP auctions in the single-item case (notice that this means SCP implies the characterization, and we do not need to worry about Global-SCP, which we did not define). The characterization is as follows: For some reserve price $r$, the auction confirms the item to the highest bidder above $r$, and burns $r$. The payment function may be determined arbitrarily (as long as it respects individual rationality and burn-balance).
@@
We prove this by De-Morgan’s law, i.e., we show that if $2$-SCP is violated, then so is $1$-SCP. Consider a beneficial $2$-SCP collusion $\Omega$, going from setting $A$ to $B$. By Lemma 3.9, we can assume there is a single bidder $i$ who changes their bid. If the other bidder in the collusion $j$ does not win the item under $B$, then their utility in $B$ is 0, and by individual rationality of $A$, they do not gain from the collusion. Thus, the $1$-SC that only involves the miner and agent $i$ is beneficial as well.
@@
In (Chung and Shi, 2023), they show an impossibility for UIC + $1$-SCP mechanisms with a finite block size. However, with an infinite block size, (Chung et al., 2024) present the discount auction, which is UIC+$1$-SCP. We complement these results by characterizing UIC+$2$-SCP mechanisms in the infinite block-size case:
@@ 5. Discussion
Going beyond our discussion of collusion size, there are other aspects of the collusion that may still be hard. In terms of computational complexity, it may be hard to identify how to implement a successful collusion. This depends on the computational model of how the problem is represented. Importantly, at least at surface-level, it seems like our result does not change the basic complexity. Consider two natural choices. If we assume that the problem is given as a full discrete table mapping all possible types (values) to outcomes (confirmation, payment, and burn), then the problem of finding a collusion is a matter of iterating over all possible types, and finding two settings where a beneficial collusion exists. This is polynomial (quadratic) in the input size, and true even without our reduction. On the other hand, if one assumes as in (Golowich and Li, 2022) that the problem is given as a Boolean circuit, then the problem of finding a beneficial $2$-SCP is coNP-complete.
@@
In terms of communication complexity, it is important to understand how many rounds of communication are needed between the colluders to establish the collusion, given that the values of the bidders are private information. This is especially important when the collusion coordination must be very quick, as Ethereum mines a block every $12$ seconds on average (Zhang et al., 2021). Ideally (for the colluders), no communication would be needed, and a tacit collusion strategy exists that each may apply on their own. It is clear that at most $2$ additional rounds of communication are required for a collusion: One round where all bidders share their values with the miner, another one where the miner tells each bidder what value to bid, and then the mandatory round where all bidders make their bids. It is interesting to understand in what settings the first two rounds can be avoided. Moreover, this suggestion leaves all the bargaining power in the hands of the miner, and this may not be acceptable for the bidders. It is interesting to know the “price of bargaining” in terms of communication, if one was to instead follow an established process of bargaining (Rubinstein, 1982; Compte and Jehiel, 2010).
@@
  • A separation of $1$-SCP and $2$-SCP with a finite block size greater than 11. The discount auction (Chung et al., 2024) shows a separation in the infinite block size case. Our Lemma 4.2 shows they are the same with a block size of $1$. What happens in-between?
  • A Global-SCP mechanism which is $1$-SCP, but not $2$-SCP. This is an open question of (Gafni and Yaish, 2025) that remains open. Global-SCP is weaker than SCP, so this is not ruled out by our results.
@@ Appendix A Missing Proofs
(Fully Burned Posted-prices $\implies$ UIC + $2$-SCP) Posted-prices are UIC as they satisfy monotonicity and critical bid (Myerson, 1981). Burning the payments does not affect bidder incentives, and so this is true also when the payments are burned. The miner utility is always zero (since all payments are burned), and so the value of any coalition is the utility of its bidders. For any coalition $C$ of bidders, their utility is maximized when each of them is confirmed according to $1[b_{i}\geq r]$, and so it is maximized in the original mechanism without any collusion.
@@ Appendix B A Separation of Active-Miner / Passive-Miner Models
Note that this is not the same as (Chung and Shi, 2023)’s “Burning Second-Price Auction”, despite the similarity in the name.
@@
The fully-burned second-price auction is UIC (as burns are irrelevant to UIC, and in terms of confirmation and payment it is equivalent to the single-item second-price auction, which is UIC). (Chung and Shi, 2023) show there is no UIC+$1$-SCP mechanism for a finite number of items, where miners can be active. In particular, consider if the miner and the winning bidder collude so that the miner drops the second-highest bid, thus increasing the winner’s utility (lower payment), but not affecting the zero miner’s utility. ∎
@@ Appendix C Theory and Evidence of Collusion in TFMs and Auctions
By now, there is overwhelming evidence of miners engaging in strategic behavior. Miners strategically withhold transactions to increase fees in Bitcoin (Lehar and Parlour, 2020), and follow a pattern of colluding with each other in doing so. (Li et al., 2024) finds mining cartels in Monacoin engage in what is known as a Selfish Mining attack (Eyal and Sirer, 2018). Miners extract value through MEV (Daian et al., 2019; Weintraub et al., 2022), and miners use intricacies of the time-stamping protocol to increase revenue (Yaish et al., 2023). (Das et al., 2022) find evidence for shill-bidding used to increase miner revenue in NFT sales. They estimate the revenue generated in this way to be at least $13$ million USD. This directly relates to the SCP notion, as the form of shill bidding they consider falls under the definition of SCP (a colluding bidder issues bids with no intention of winning, to increase the payment to the miner in an English auction). They also find evidence for bid shielding, which is more of a behavioral pattern: In an auction where bidders can retract their bids, a high bidder deters competition, and at the last minute pulls out so that a low bidder can win at a low price. The shielded bids amount to a total of about $940,000$ USD. While this is a form of user-user collusion, it is interesting that it is an order of magnitude less manifested, somewhat supporting the SCP’s notion focus on miner-user collusion. However, (Gehrlein et al., 2025) observes user-user collusion in the Polkadot Parachains candle auctions, and show evidence that bidders engage in turn-taking, so as to jointly lower payments (‘you avoid bidding against me, and I will avoid bidding against you’).
@@
There is extensive evidence document collusion in traditional auctions. First, there is the legal evidence: At some points (in particular, this was prevalent in the 1980s), more than $50\%$ of Antitrust cases in the US involved bid rigging in auctions (Porter and Zona, 1993), and this remained a substantial issue thereafter (Krishna, 2010). Second, there is the industry insiders / anecdotal evidence: (Graham and Marshall, 1987) mention an auctioneer who said “that in 40 years of auctioneering, he had yet to attend an auction at which a ring was not present”. Third, there is an extensive economic literature developing methods to identify collusion empirically through statistical tests, based on publicly known data. This approach relies on identifying “tells” that would arise in collusive bidding but not in competitive bidding. In user-user collusion, bidders may avoid undercutting one another (Baránek et al., 2025; Chassang et al., 2022). Winning bids will be drawn from a distribution that corresponds to and makes sense as real valuations, but losing bids will not be constrained in this way (Porter and Zona, 1993). Fourth, controlled lab experiments show that bidders that communicate and make transfers significantly reduce miner revenue, which is interpreted as facilitating user-user collusion (Agranov and Yariv, 2018).
@@
Intuitively, there are several issues that prohibit a larger-size collusion. (Stigler, 1964) raises the issue of the possible lack of trust among the colluders. This may prohibit the collusion in at least two ways: (1) A colluder may implement a policy that violates the agreed-upon collusion. (Ferreira et al., 2021) explores this issue in the blockchain setting, by characterizing TFMs robust only to a collusion where it is incentive-compatible for bidders to follow the collusion, and not further deviate from it for self-gain. To counter that, (Kelkar et al., 2025) introduce the idea of smart collusion, where colluders can impose penalties on defectors, in line with the models of many works in traditional economic theory.
@@
(2) A colluder may defect and report the existence of the collusion. (Schrepel, 2019) writes that the “leniency program”, which allows members of cartels to self-report, ”has become the most effective tool in the fight against cartels”. Another issue in a larger-size collusion is that of bargaining. Results in cooperative game theory (Compte and Jehiel, 2010) show that bargaining among a larger coalition results in higher loss of economic efficiency.
@@
Empirical evidence shows that a small-sized collusion is more prevalent than a larger-size one, across different works (Posner, 1970; Hay and Kelley, 1974; Frass and Greer, 1977).

### v2

@@ 4. Corollaries Of The Main Characterization
Consider any $2$-SCP auction $a,p,\beta$. By our characterization, it is also SCP. (welfareIncreasingCollusion) show that SCP implies Global-SCP, and provide a characterization of Global-SCP auctions in the single-item case (notice that this means SCP implies the characterization, and we do not need to worry about Global-SCP, which we did not define). The characterization is as follows: For some reserve price $r$, the auction confirms the item to the highest bidder above $r$, and burns $r$. The payment function may be determined arbitrarily (as long as it respects individual rationality and burn-balance).
@@
We prove the contrapositive, i.e., we show that if $2$-SCP is violated, then so is $1$-SCP. Consider a beneficial $2$-SCP collusion $\Omega$, going from setting $A$ to $B$. By Lemma 3.10, we can assume there is a single bidder $i$ who changes their bid. If the other bidder in the collusion $j$ does not win the item under $B$, then their utility in $B$ is 0, and by individual rationality of $A$, they do not gain from the collusion. Thus, the $1$-SC that only involves the miner and agent $i$ is beneficial as well.
@@
In (chung2023foundations), they show an impossibility for UIC + $1$-SCP mechanisms with a finite block size. However, with an infinite block size, (chung2024collusion) present the discount auction, which is UIC+$1$-SCP. We complement these results by characterizing UIC+$2$-SCP mechanisms in the infinite block-size case:
@@ 5. Discussion
Going beyond our discussion of collusion size, there are other aspects of the collusion that may still be hard. In terms of computational complexity, it may be hard to identify how to implement a successful collusion. This depends on the computational model of how the problem is represented. Importantly, at least at surface-level, it seems like our result does not change the basic complexity. Consider two natural choices. If we assume that the problem is given as a full discrete table mapping all possible types (values) to outcomes (confirmation, payment, and burn), then the problem of finding a collusion is a matter of iterating over all possible types, and finding two settings where a beneficial collusion exists. This is polynomial (quadratic) in the input size, and true even without our reduction. On the other hand, if one assumes as in (golowichLi2022) that the problem is given as a Boolean circuit, then the problem of finding a beneficial $2$-SCP is coNP-complete.
@@
In terms of communication complexity, it is important to understand how many rounds of communication are needed between the colluders to establish the collusion, given that the values of the bidders are private information. This is especially important when the collusion coordination must be very quick, as Ethereum mines a block every $12$ seconds on average (ethereumLatency). Ideally (for the colluders), no communication would be needed, and a tacit collusion strategy exists that each may apply on their own. It is clear that at most $2$ additional rounds of communication are required for a collusion: One round where all bidders share their values with the miner, another one where the miner tells each bidder what value to bid, and then the mandatory round where all bidders make their bids. It is interesting to understand in what settings the first two rounds can be avoided. Moreover, this suggestion leaves all the bargaining power in the hands of the miner, and this may not be acceptable for the bidders. It is interesting to know the “price of bargaining” in terms of communication, if one was to instead follow an established process of bargaining (rubinsteinBargaining; compteJehiel).
@@
  • A separation of $1$-SCP and $2$-SCP with a finite block size greater than $1$. The discount auction (chung2024collusion) shows a separation in the infinite block size case. Our Lemma 4.2 shows they are the same with a block size of $1$. What happens in-between?
  • A Global-SCP mechanism which is $1$-SCP, but not $2$-SCP. This is an open question of (welfareIncreasingCollusion) that remains open. Global-SCP is weaker than SCP, so this is not ruled out by our results.
@@ Appendix A Missing Proofs
(Fully Burned Posted-prices $\implies$ UIC + $2$-SCP) Posted-prices are UIC as they satisfy monotonicity and critical bid (myerson1981optimal). Burning the payments does not affect bidder incentives, and so this is true also when the payments are burned. The miner utility is always zero (since all payments are burned), and so the value of any coalition is the utility of its bidders. For any coalition $C$ of bidders, their utility is maximized when each of them is confirmed according to $1[b_{i}\geq r]$, and so it is maximized in the original mechanism without any collusion.
@@ Appendix B A Separation of Active-Miner / Passive-Miner Models
Note that this is not the same as (chung2023foundations)’s “Burning Second-Price Auction”, despite the similarity in the name.
@@
The fully-burned second-price auction is UIC (as burns are irrelevant to UIC, and in terms of confirmation and payment it is equivalent to the single-item second-price auction, which is UIC). (chung2023foundations) show there is no UIC+$1$-SCP mechanism for a finite number of items, where miners can be active. In particular, consider if the miner and the winning bidder collude so that the miner drops the second-highest bid, thus increasing the winner’s utility (lower payment), but not affecting the zero miner’s utility. ∎
@@ Appendix C Theory and Evidence of Collusion in TFMs and Auctions
By now, there is overwhelming evidence of miners engaging in strategic behavior. Miners strategically withhold transactions to increase fees in Bitcoin (minersCollusionParlour), and follow a pattern of colluding with each other in doing so. (statisticalDetectionSelfishMining) finds mining cartels in Monacoin engage in what is known as a Selfish Mining attack (selfishMining). Miners extract value through MEV (flashboys2; flashbots2022flashbots), and miners use intricacies of the time-stamping protocol to increase revenue (yaish2023uncle). (NFTSecurity) find evidence for shill-bidding used to increase miner revenue in NFT sales. They estimate the revenue generated in this way to be at least $13$ million USD. This directly relates to the SCP notion, as the form of shill bidding they consider falls under the definition of SCP (a colluding bidder issues bids with no intention of winning, to increase the payment to the miner in an English auction). They also find evidence for bid shielding, which is more of a behavioral pattern: In an auction where bidders can retract their bids, a high bidder deters competition, and at the last minute pulls out so that a low bidder can win at a low price. The shielded bids amount to a total of about $940,000$ USD. While this is a form of user-user collusion, it is interesting that it is an order of magnitude less manifested, somewhat supporting the SCP’s notion focus on miner-user collusion. However, (candleAuctionsPolkadot) observes user-user collusion in the Polkadot Parachains candle auctions, and show evidence that bidders engage in turn-taking, so as to jointly lower payments (‘you avoid bidding against me, and I will avoid bidding against you’).
@@
There is extensive evidence document collusion in traditional auctions. First, there is the legal evidence: At some points (in particular, this was prevalent in the 1980s), more than $50\%$ of Antitrust cases in the US involved bid rigging in auctions (porterZona1993), and this remained a substantial issue thereafter (krishna2002). Second, there is the industry insiders / anecdotal evidence: (grahamMarshall) mention an auctioneer who said “that in 40 years of auctioneering, he had yet to attend an auction at which a ring was not present”. Third, there is an extensive economic literature developing methods to identify collusion empirically through statistical tests, based on publicly known data. This approach relies on identifying “tells” that would arise in collusive bidding but not in competitive bidding. In user-user collusion, bidders may avoid undercutting one another (musolf2025; kawaiNa). Winning bids will be drawn from a distribution that corresponds to and makes sense as real valuations, but losing bids will not be constrained in this way (porterZona1993). Fourth, controlled lab experiments show that bidders that communicate and make transfers significantly reduce miner revenue, which is interpreted as facilitating user-user collusion (agranovYariv).
@@
Intuitively, there are several issues that prohibit a larger-size collusion. (Stigler1964) raises the issue of the possible lack of trust among the colluders. This may prohibit the collusion in at least two ways: (1) A colluder may implement a policy that violates the agreed-upon collusion. (ferreira2021dynamic) explores this issue in the blockchain setting, by characterizing TFMs robust only to a collusion where it is incentive-compatible for bidders to follow the collusion, and not further deviate from it for self-gain. To counter that, (smartCollusion) introduce the idea of smart collusion, where colluders can impose penalties on defectors, in line with the models of many works in traditional economic theory.
@@
(2) A colluder may defect and report the existence of the collusion. (harvardLawCartels) writes that the “leniency program”, which allows members of cartels to self-report, ”has become the most effective tool in the fight against cartels”. Another issue in a larger-size collusion is that of bargaining. Results in cooperative game theory (compteJehiel) show that bargaining among a larger coalition results in higher loss of economic efficiency.
@@
Empirical evidence shows that a small-sized collusion is more prevalent than a larger-size one, across different works (posner70; hayKelley74; FrassGreer77).